A screen is placed a distance L from a single slit of width a, which is illuminated with light of wavelength ?. Assume L
Answer #1

the formula -

? = ws/D

Rearrange to get in terms of s(slit spacing)

s = ?D/w

Throw in your values (remember to convert to m)
Answer #2

The general formula to apply here is y/L = m(wavelength/a

We need delta y alone by subtracting, so...

y - y = (m2)(wavelength)L/a - (m1?)(wavelength)L/a

Since L, a and wavelength are not cocnstant, we get

(delta y) = (m2-m1)(wavelength)L/a

Solving for a gives us

a = (m2-m1)(wavelength)L/(delta y)

Rewritten, that is...

a=\frac{\left (m_{2}-m_{1})L\lambda }{\Delta y}